# Identifying Challenges in Designing, Developing and Evaluating Data Visualizations for Large Displays

### v1

@@ 2.1 Limited Resources for Building LD Infrastructure
Building LD infrastructure requires a large space to enable users to freely move and interact. This need imposed constraints in building and working with LDs. One participant (P1) mentioned “Usually we moved all the chairs and tables because my research was focusing on making people like [minimum] three people to collaborate in their room and use the large display in an environment so they can explore the data ”. In addition to space allocation, we need to implement effective cooling systems to manage the heat generated during LD operations. Prolonged use of LDs, especially in environments with inadequate ventilation or cooling, can lead to increased heat buildup, which could result in an uncomfortable working environment. To address these issues, we need to allocate sufficient space and install proper ventilation systems, which can be time-consuming and costly. As a result, this infrastructure typically involves fixed configurations that cannot be easily relocated.
@@ 2.2 Scaling Data Visualization Designs for LDs
To design a data visualization prototype, we often start by sketching the designs. LDs often have different sizes, resolutions, and aspect ratios compared to paper or surfaces typically used for sketching. This disparity made it challenging for participants to sketch their data visualizations for LDs. Participants highlighted the need for flip-charts to create a fake wall, catering to the specific characteristics of LDs, in order to create sketches of their design on it. P3 said “what we did, which would just get flipchart paper, which, like one and a half flipchart paper, kind of corresponded to what we had in the screen wise for one screen, right? So you could then essentially tape, so we could take the whole wall with it and we had the wall that was big enough to actually simulate [the LD] …So later, we had a wall and like 4 flipchart stands as well so that you can do the curve thing as well. It looked pretty silly, but it was actually very effective”.
@@ 2.3 Limited Development Resources and Tools for LDs
In addition to all of these challenges, available tools, software, and toolkits specifically tailored for LD prototyping is limited. While there may be some existing solutions, their functionalities and capabilities may not fully meet the requirements and challenges posed by LDs. Due to the limited availability of support and resources, developers often need to adapt or modify existing tools or create custom solutions to meet the requirements of LD prototyping. P4 said “I didn’t get too much help from …just in that because this is like a very specific platform that I needed”.
@@ 2.4 Evaluating Data Visualizations on LDs with People
Evaluating data visualizations on LDs with people can be challenging as people often are unfamiliar with LD technology, as it is not widely accessible or commonly used. P3 said “ We grew up with this thing that you don’t want to break the TV. It’s expensive and I kind of feel that this is kind of cultural thing that people are afraid to [touch] them, right? and I think this is a challenge now [and it will not be a challenge] 20 years in the future”.

### v3

@@ 2.1 Limited Resources for Building LD Infrastructure
Building LD infrastructure requires a large space to enable users to freely move and interact. This need imposed constraints in building and working with LDs. One participant (P1) mentioned “We moved all the chairs and tables because my research was focusing on making [minimum] three people to collaborate and use the large display so they can explore the data.” In addition to space allocation, we need to implement effective cooling systems to manage the heat generated during LD operations. Prolonged use of LDs, especially in environments with inadequate ventilation or cooling, can lead to increased heat buildup, which could result in an uncomfortable working environment. To address these issues, we need to allocate sufficient space and install proper ventilation systems, which can be time-consuming and costly. As a result, this infrastructure typically involves fixed configurations that cannot be easily relocated.
@@ 2.2 Scaling Data Visualization Designs for LDs
To design a data visualization prototype, we often start by sketching the designs. LDs often have different sizes, resolutions, and aspect ratios compared to paper or surfaces typically used for sketching. This disparity made it challenging for participants to sketch their data visualizations for LDs. Participants highlighted the need for flip-charts to create a fake wall, catering to the specific characteristics of LDs, in order to create sketches of their design on it. P3 said “what we did, which would just get flipchart paper, kind of corresponded to what we had in the screen wise for one screen. So we could take the whole wall with it and we had the wall that was big enough to actually simulate [the LD] …So later, we had a wall and like 4 flipchart stands as well so that you can do the curve thing as well. It was actually very effective.”
@@ 2.3 Limited Development Resources and Tools for LDs
In addition to all of these challenges, available tools, software, and toolkits specifically tailored for LD prototyping is limited. While there may be some existing solutions, their functionalities and capabilities may not fully meet the requirements and challenges posed by LDs. Due to the limited availability of support and resources, developers often need to adapt or modify existing tools or create custom solutions to meet the requirements of LD prototyping. P4 said “I didn’t get too much help from …just in that because this is like a very specific platform that I needed.”
@@ 2.4 Evaluating Data Visualizations on LDs with People
Evaluating data visualizations on LDs with people can be challenging as people often are unfamiliar with LD technology, as it is not widely accessible or commonly used. P3 said “We grew up with this thing that you don’t want to break the TV. It’s expensive and people are afraid to [touch] them. I think this is a challenge now [and it will not be a challenge] 20 years in the future.”